The sum of Mr. Micklow and Mrs. Craft's age is 55. The difference of their age is 3. What are their ages?
Found 2 solutions by ikleyn, mananth:
Answer by ikleyn(52787)

.

   Mrs. +  Mr.  = 55

    x   + (x+3) = 55

      2x        = 55 - 3

      2x        = 52

       x        = 52/2 = 26.


Mrs. is 26 years old.  Mr. is 26 + 3 = 29 years old.


CHECK.  26 + 29 = 55 years, in total.   ! Correct !

Solved.



Answer by mananth(16946)

let Mr. Micklow's age be x years
and Mrs. Craft's age be y years
The sum of Mr. Micklow and Mrs. Craft's age is 55.
So x+y =55
The difference of their age is 3. What are their ages?
x-y =3
Add the two equations
we get 2x = 58
x= 58/2
x=29
substitute x
in x+y =54
29+y =55
y= 26
Mr. Micklow's age 29 years
and Mrs. Craft's age 26 years